# A Review of Scalable and Privacy-Preserving Multi-Agent Frameworks for Distributed Energy Resource Control

### v1

@@ IV-A4 Multi-agent reinforcement learning
Remark 1
IV-B Privacy Leakages
To ensure the warranted use of private information from all stakeholders, it is crucial to synthesize privacy preservation techniques into the design of scalable DER control strategies. Toward this goal, we summarize the typical adversaries in multi-agent computing frameworks, including external eavesdroppers, honest-but-curious agents, and the SO and/or coordinators/aggregators, each representing a distinct type of threat with different attack vectors. By examining these three adversaries, we cover a broad spectrum of external, internal, and hierarchical privacy risks, offering a comprehensive understanding of the threats in multi-agent systems. This helps guide the design of privacy-preserving multi-agent frameworks that require privacy guarantees in different operational scenarios.
IV-B1 External eavesdroppers
IV-B2 Honest-but-curious agents
IV-B3 The system operator and/or coordinators/aggregators
@@
V Privacy Preservation Techniques
TABLE II: Differential-Privacy (DP)-based Scalable and Privacy-Preserving Methods.
Method Reference Problem Structure Adversaries Key Features
@@ V Privacy Preservation Techniques
DP [175] Consensus for multi-agent system Distributed Byzantine, malicious agents, $\epsilon$-DP 1-A subset of agents is adversarial; 2-achieve resilient asymptotic consensus with correctness, accuracy and DP properties.
@@
           DP: Differential privacy
In this section, we explore the details of both mainstream and emerging privacy preservation techniques for scalable multi-agent frameworks and demonstrate their applications in DER control problems (see Tables II, III, IV,V). Furthermore, we discuss potential research and development directions for each privacy preservation technique.
V-A Differential Privacy
The concept of differential privacy (DP), first introduced by Dwork [188, 189], captures the increased risk to one’s privacy incurred by participating in a database. By adding random noises to the database, a curator (or SO) can release statistical information output of a data analysis result without compromising any individuals’ privacy. Owing to its rigorous mathematical definition, DP has been a de facto standard in developing privacy reservation techniques. DP-based methods can quantify the privacy loss at a differential change in a database (i.e., adding or removing one entry), described by a privacy parameter $\epsilon$ that captures the privacy loss. DP ensures that privacy is preserved regardless of the combination of computations performed on the dataset, providing strong privacy guarantees against arbitrary adversaries, e.g., any re-identification attack [190]. DP is a powerful privacy preservation architecture to make confidential data widely available for data analysis in the broad areas of artificial intelligence [191, 192, 193], power and energy systems [194, 195, 196], and the internet of things [197, 198]. To aid in understanding, the definition of $\epsilon$-DP is given here [189]:
The rigorous theory foundation of DP has led to extensive privacy preservation applications in multi-agent systems and DER control in power systems. By adding well-calibrated noises into the computation process, DP-based methods can obscure the attributes of any single individual’s data (e.g., smart meter readings) without affecting grid-level and/or DER-level objectives. Based on DP, Hale and Egerstedt [174] develop a privacy-preserving primal-dual optimization framework for multi-agent convex programs and solve it using the PGD. It keeps each agent’s state trajectory private from all other agents and any external eavesdroppers. Han et al. [177] develop a distributed privacy-preserving optimization algorithm based on DP to preserve the privacy of the participating agents in constrained optimizations. To broaden the range of adversaries, Fiore and Russo [175] design a DP-based consensus algorithm for multi-agent systems where a subset of agents could be honest-but-curious. In [55], a DP-based privacy preservation algorithm is developed to protect consumers’ smart meter data. Dvorkin et al. [56] develop an adversarial inference model based on DP that first questions the privacy properties of distributed OPF. Subsequently, the authors develop a differentially private variant of the ADMM to ensure information privacy during information exchanges between neighbors. This model is later extended in [176] for the distributed optimization of AC power flow problems. In [57], a DP-based aggregation algorithm is proposed to compensate for solar power fluctuations and protect customers’ personal information. In [194], a DP-based obfuscation mechanism is proposed with guarantees of AC feasibility to protect the private parameters of transmission lines and transformers.
To summarize, DP has become a golden rule in the domain of privacy preservation, rapidly evolving alongside the advancement of scalable multi-agent frameworks. The potential of DP can be further explored from the following directions: (1) Reduce the privacy-accuracy gap. DP-based methods commonly suffer from loss of accuracy caused by the added noise. Research shows that a balance between privacy and accuracy can be achieved via the design of carefully calibrated noises [199, 200]. (2) Extension of DP for both privacy (passive adversaries) and security (active adversaries) scenarios. When faulty agents maliciously deviate from the computing policy or network communication protocol, the effectiveness of DP in maintaining both privacy and security can be compromised [201, 202, 203]. The co-design of a privacy-preserving and cybersecure multi-agent framework is worth further investigation. (3) Enhanced compatibility for the next generation of learning-aided methods. DP has shown strong cohesion in preserving privacy for learning-aided methods, including training train neural networks for deep learning models [204], employing stochastic gradient descent for machine learning [205], reducing sample complexity with new expansion on DP [206]. The rapid evolution of DP also shows strong compatibility in addressing emerging privacy concerns in learning-aided approaches, such as the deployment of large language models in the electric power sector [207].
V-B Cryptographic Methods
@@
TABLE III: Encryption-Decryption (ED)-based Scalable and Privacy-Preserving Methods.
Method Reference Problem Structure Adversaries Key Features
@@
           Refers to active adversaries    ED: Encryption-decryption
V-B1 Encryption-decryption-based methods
Encryption-decryption (ED)-based methods utilize a cryptosystem that typically consists of three components: An encryption algorithm, a decryption algorithm, and key management. Specifically, a plaintext $m$ is encrypted into a ciphertext $\mathcal{E}(m)$ using an encryption function $\mathcal{E}(\cdot)$. By applying a decryption function $\mathcal{D}(\cdot)$ to the ciphertext, the original plaintext can be correctly retrieved as $m=\mathcal{D}(\mathcal{E}(m))$. Fig. 5 shows the realization of secure communications using a cryptosystem. A sender sends some sensitive plaintexts to a receiver in the form of ciphertexts using a cryptosystem such that any party intercepting/eavesdropping on the communication channel only has access to the ciphertexts, instead of knowing the plaintexts.
Among various cryptosystems, homomorphic cryptosystems are well-suited for multi-agent computing and communications. Essentially, a homomorphic cryptosystem enables users to perform computations on encrypted data without having to decrypt it first. The homomorphic properties are typically necessary for performing secure arithmetic operations in multi-agent systems [10, 208, 209, 61, 62, 219]. Homomorphic schemes can be classified according to the types of mathematical operations that can be performed on ciphertexts: 1) Partially homomorphic that supports either addition or multiplication operation, but not both simultaneously, and 2) fully homomorphic that concurrently support addition and multiplication operations. For a cryptosystem to be fully homomorphic, it needs to satisfy:
Consider the PGD in Section IV-A3, we give an example of integrating ED into PGD-based scalable multi-agent frameworks. As shown in Fig. 6, plaintexts are calculated and transmitted directly without privacy protection between agents and the system (coordinator) in a primal-dual-based computing scheme. By using ED, agents can encrypt private information (e.g., decision variables, private coefficients, subgradients, objective functions [10, 208]) and then communicate with the coordinator only via ciphertexts. The coordinator can access, aggregate, and compute ciphertexts based on the cryptosystem’s homomorphic properties. The primal and dual updates can be executed in the space of ciphertexts.
In the power system field, ED-based methods are compatible with power systems’ complex computing and communication structures for transmitting sensitive data, such as customer load profiles, operational status, and control commands. Moreover, they are integrable with electric engineering standards including IEC 62351, IEEE 1815-2012 (DNP3), and NERC reliability standards [224, 225, 226]. To preserve the private voltage and current measurements, Wu et al. [62] develop a privacy-preserving distributed OPF algorithm based on partially homomorphic cryptosystems. To eliminate the privacy concerns of economic dispatch problems in microgrids, a homomorphically encrypted algorithm is developed to achieve consensus without disclosing agents’ private or sensitive state information [227]. He et al. [210] develop a computationally efficient data aggregation scheme based on public key cryptography to prevent the extraction of consumers’ electricity consumption information against internal and external attackers.
@@ V-B1 Encryption-decryption-based methods
TABLE IV: Secret Sharing (SS)-based Scalable and Privacy-Preserving Methods.
Method Reference Problem Structure Adversaries Key Features
@@
           Refers to active adversaries    SS: Secret sharing
V-B2 Secret sharing-based methods
Secret sharing (SS) is a lightweight cryptographic protocol that can split a secret into multiple shares and distribute the shares among a group of participants. The essential idea behind SS is to ensure that the secret can only be reconstructed by combining an adequate number of shares. Meanwhile, any subset of shares smaller than a threshold yields no useful information about the secret. The procedures of SS, including the division of shares and the reconstruction of secrets, are given in Fig. 7.
Figure 7: A depiction of SS on the division of shares and the secret reconstruction process.
Shamir’s SS [234] is a well-known SS scheme in which the secret shares are generated using a polynomial. Specifically, Shamir’s SS is developed based on the concept of polynomial interpolation, defined as [235]:
Theorem 1 (Polynomial interpolation). Let $\{(\varsigma_{1},y_{1}),\ldots,$ $(\varsigma_{\bar{d}},y_{\bar{d}})\}$ $\subseteq\mathbb{R}^{2}$ be a set of points whose values of $\varsigma_{d}$ are all distinct. Then, there exists a unique polynomial $\mathcal{Y}$ of degree $\bar{d}-1$ that satisfies $y_{d}=\mathcal{Y}(\varsigma_{d}),\forall d=1,\ldots,\bar{d}$. $\blacksquare$
Theorem 1 states that a minimum number of $d+1$ points equal to the degree of the polynomial are required to reconstruct the secret. This ensures information-theoretic security, meaning that even if an adversary obtains some shares, it is impossible to reconstruct the secret unless they have acquired the quorum number of shares. SS is highly confidential and has a variety of privacy preservation applications in multi-agent systems, including the distribution of cryptographic keys, the management of access control in distributed systems [236], and the protection of sensitive data [237, 11]. For example, an SS-based algorithm is developed in [11] to solve the consensus problem and protect each individual’s private information. In [237], a novel cloud storage system is proposed based on SS to protect sensitive electronic health records.
SS has also demonstrated strong potential in preserving privacy for DER control and other power system applications. Adopting SS, Nabil et al. [238] design a privacy-preserving detection scheme to identify electricity theft from malicious consumers. Only masked meter readings from consumers are collected and sent to the SO, ensuring privacy protection and preventing data leakage. In [63], an SS-based cooperative EV charging control protocol is developed to achieve overnight valley filling without compromising the privacy of EV owners’ charging profiles. The distributed protocol enjoys high computation efficiency and accuracy. In [232], a privacy-preserving communication protocol based on SS is proposed for vehicle-to-grid integration. The proposed protocol ensures that the existing battery charge level, the quantity of replenished energy, and the duration of EVs being plugged in remain undisclosed to aggregators. In summary, as a threshold scheme based on polynomials and finite geometries, Shamir’s SS is well-suited for secure computation and key sharing in cryptographic applications among multiple stakeholders. Some potential future directions for SS-based methods: (1) Homomorphic secret sharing. Combining SS with homomorphic properties allows computations to be performed on the shared data without revealing the secret. It is worth investigating efficient homomorphic operations within the SS framework for privacy-preserving computations in scalable multi-agent systems. (2) Threshold cryptography in dynamic environments. Traditional SS requires a predefined number of participants. However, real-world power system applications often involve dynamic groups and changing environments. Research could focus on adapting SS to handle dynamic groups, where agents can join or leave without compromising the shared secret. (3) Application-specific adaptations. The research could focus on how SS can be adapted to protect sensitive data in a broader line of applications, such as consensus mechanisms, blockchain systems, and distributed storage systems.
TABLE V: Miscellaneous and Emerging Scalable and Privacy-Preserving Methods.
Method Reference Problem Structure Adversaries Key Features

### v2

@@ 3.2.3 The system operator and/or coordinators/aggregators
4 Privacy Preservation Techniques
In this section, we explore the details of both mainstream and emerging privacy preservation techniques for scalable multi-agent frameworks and demonstrate their applications in DER control problems (see Tables LABEL:table_DP_review, LABEL:table_ED_literature, LABEL:table_SS_literature,LABEL:table_others_literature). Furthermore, we discuss potential research and development directions for each privacy preservation technique.
4.1 Differential Privacy
The concept of differential privacy (DP), first introduced by Dwork [183, 184], captures the increased risk to one’s privacy incurred by participating in a database. By adding random noises to the database, a curator (or SO) can release statistical information output of a data analysis result without compromising any individuals’ privacy. Owing to its rigorous mathematical definition, DP has been a de facto standard in developing privacy preservation/protection techniques (Note that ‘preservation’ can imply a stronger standard than ‘protection’. For example, privacy is not completely preserved by DP since it only limits privacy leakage and does not eliminate it. In this paper, we refer to the broad concept of ‘protection’ when using both terms). DP-based methods can quantify the privacy loss at a differential change in a database (i.e., adding or removing one entry), described by a privacy parameter $\epsilon$ that captures the privacy loss. DP ensures that privacy is preserved regardless of the combination of computations performed on the dataset, providing strong privacy guarantees against arbitrary adversaries, e.g., any re-identification attack [185]. DP is a powerful privacy preservation architecture to make confidential data widely available for data analysis in the broad areas of artificial intelligence [186, 187, 188], power and energy systems [189, 190, 191], and the internet of things [192, 193]. To aid in understanding, the definition of $\epsilon$-DP is given here [184]:
The rigorous theoretical foundation of DP has led to extensive privacy preservation applications in multi-agent systems and DER control in power systems. For example, DP-based methods can add calibrated noise into smart meter data and/or into the computing process, therefore protecting the attributes of any single individual’s smart meter readings without affecting grid-level and/or DER-level objectives. DP-based methods have been integrated into scalable methods to achieve privacy protection for power system applications, such as protecting consumers’ smart meter data [73], power flow problems [74, 194], and EV charging control [195]. Based on DP, Hale and Egerstedt [196] develop a privacy-preserving primal-dual optimization framework for multi-agent convex programs and solve it using the PGD. It keeps each agent’s state trajectory private from all other agents and any external eavesdroppers. Han et al. [197] develop a distributed privacy-preserving optimization algorithm based on DP to preserve the privacy of the participating agents in constrained optimizations. To broaden the range of adversaries, Fiore and Russo [198] design a DP-based consensus algorithm for multi-agent systems where a subset of agents could be honest-but-curious. In [73], a DP-based privacy preservation algorithm is developed to protect consumers’ smart meter data. Dvorkin et al. [74] develop an adversarial inference model based on DP that first questions the privacy properties of distributed OPF. Subsequently, the authors develop a differentially private variant of the ADMM to ensure information privacy during information exchanges between neighbors. This model is later extended in [194] for the distributed optimization of AC power flow problems. In [75], a DP-based aggregation algorithm is proposed to compensate for solar power fluctuations and protect customers’ personal information. In [189], a DP-based obfuscation mechanism is proposed with guarantees of AC feasibility to protect the private parameters of transmission lines and transformers. Lee and Choi [195] develop a DP-based multilevel deep RL algorithm for privacy-preserving EV charging operations, ensuring optimized data privacy, revenue, and energy costs.
To summarize, DP has become a foundational principle in the privacy protection field, with wide-ranging power system applications, and is evolving rapidly alongside advancements in developing scalable multi-agent frameworks. The potential of DP can be further explored from the following directions: (1) Reduce the privacy-accuracy gap. DP-based methods commonly suffer from loss of accuracy caused by the added noise. Research shows that a balance between privacy and accuracy can be achieved via the design of carefully calibrated noises [199, 200]. (2) Extension of DP for both privacy (passive adversaries) and security (active adversaries) scenarios. When faulty agents maliciously deviate from the computing policy or network communication protocol, the effectiveness of DP in maintaining both privacy and security can be compromised [201, 202, 203]. The co-design of a privacy-preserving and cybersecure multi-agent framework is worth further investigation. (3) Enhanced compatibility for the next generation of learning-aided methods. DP has shown strong cohesion in preserving privacy for learning-aided methods, including training train neural networks for deep learning models [204], employing stochastic gradient descent for machine learning [205], reducing sample complexity with new expansion on DP [206]. The rapid evolution of DP also shows strong compatibility in addressing emerging privacy concerns in learning-aided approaches, such as the deployment of large language models in the electric power sector [207].
Remark 3
The DP technique can be categorized into global DP and local DP depending on whether the server (e.g., SO/coordinator/aggregator) can be trusted during the computing process. In global DP, privacy is maintained by adding noise at the central server level (e.g., database query), while in local DP, privacy is preserved by individuals before sending any raw data to the central server (e.g., adding noise to smart meter data). In practice, choosing between global and local DP depends on the use case and the level of trust in the central server, i.e., global DP is common when a central server is trusted [196], while local DP is preferred when trust is an issue and individual-level privacy must be ensured [198]. Note that, the definition of centralized, distributed, and decentralized information exchange structures are categorized based on their communication patterns, i.e., how private information flows between different parties, while the categorization of DP reflects the trustworthiness of the central server. $\square$
Table 2: Differential-Privacy (DP)-based Scalable and Privacy-Preserving Methods.
Method Ref. Problem Structure Adversaries Key Features
@@ 4.1 Differential Privacy
DP [198] Consensus for multi-agent system Distributed Byzantine and malicious agents, $\epsilon$-DP 1-A subset of agents is adversarial; 2-achieve resilient asymptotic consensus with correctness, accuracy and DP properties.
@@
Table 2: (Continued)
Ref.: Reference    DP: Differential privacy
4.2 Cryptographic Methods
@@
4.2.1 Encryption-decryption-based methods
Encryption-decryption (ED)-based methods utilize a cryptosystem that typically consists of three components: An encryption algorithm, a decryption algorithm, and key management. Specifically, a plaintext $m$ is encrypted into a ciphertext $\mathcal{E}(m)$ using an encryption function $\mathcal{E}(\cdot)$. By applying a decryption function $\mathcal{D}(\cdot)$ to the ciphertext, the original plaintext can be correctly retrieved as $m=\mathcal{D}(\mathcal{E}(m))$.
Fig. 5 shows the realization of secure communications using a cryptosystem. A sender sends some sensitive plaintexts to a receiver in the form of ciphertexts using a cryptosystem such that any party intercepting/eavesdropping on the communication channel only has access to the ciphertexts, instead of knowing the plaintexts.
Among various cryptosystems, homomorphic cryptosystems are well-suited for multi-agent computing and communications. Essentially, a homomorphic cryptosystem enables users to perform computations on encrypted data without having to decrypt it first. The homomorphic properties are typically necessary for performing secure arithmetic operations in multi-agent systems [21, 217, 218, 79, 80, 219], offering significant potential for power system applications, such as optimal power flows [80], DER management [220, 221], and economic dispatch [222, 223]. Homomorphic schemes can be classified according to the types of mathematical operations that can be performed on ciphertexts: 1) Partially homomorphic that supports either addition or multiplication operation, but not both simultaneously, and 2) fully homomorphic that concurrently support addition and multiplication operations. For a cryptosystem to be fully homomorphic, it needs to satisfy:
Example 1 (Integration of ED into PGD). Consider the PGD in Section 3.1.3, we give an example of integrating ED into PGD-based scalable multi-agent frameworks.
As shown in Fig. 6, plaintexts are calculated and transmitted directly without privacy protection between agents and the system (coordinator) in a primal-dual-based computing scheme. By using ED, agents can encrypt private information (e.g., decision variables, private coefficients, subgradients, objective functions [21, 217]) and then communicate with the coordinator only via ciphertexts. The coordinator can access, aggregate, and compute ciphertexts based on the cryptosystem’s homomorphic properties. The primal and dual updates can be executed in the space of ciphertexts. $\square$
In the power systems field, ED-based methods are well-suited for real-world implementation due to their compatibility with complex computing and communication structures. This compatibility enables the secure transmission of sensitive information, including customer load profiles, power system operational status, and operator control commands. Moreover, the scalability and privacy protection capabilities of ED-based methods enhance the industry relevance of multi-agent frameworks for DER management, aligning them with electrical engineering standards such as IEC 62351, IEEE 1815-2012 (DNP3), and NERC reliability standards [39, 228, 229]. To preserve the private voltage and current measurements, Wu et al. [80] develop a privacy-preserving distributed OPF algorithm based on partially homomorphic cryptosystems. To eliminate the privacy concerns of economic dispatch problems in microgrids, a homomorphically encrypted algorithm is developed to achieve consensus without disclosing agents’ private or sensitive state information [230]. He et al. [231] develop a computationally efficient data aggregation scheme based on public key cryptography to prevent the extraction of consumers’ electricity consumption information against internal and external attackers.
Table 3: Encryption-Decryption (ED)-based Scalable and Privacy-Preserving Methods.
Method Ref. Problem Structure Adversaries Key Features
@@ 4.2.1 Encryption-decryption-based methods
Table 3: (Continued)
Refers to active adversaries    Ref.: Reference    ED: Encryption-decryption
4.2.2 Secret sharing-based methods
Secret sharing (SS) is a lightweight cryptographic protocol that can split a secret into multiple shares and distribute the shares among a group of participants. The essential idea behind SS is to ensure that the secret can only be reconstructed by combining an adequate number of shares. Meanwhile, any subset of shares smaller than a threshold yields no useful information about the secret. Shamir’s SS [236] is a well-known SS scheme in which the secret shares are generated using a polynomial. Specifically, Shamir’s SS is developed based on the concept of polynomial interpolation, defined as [237]:
Theorem 1 (Polynomial interpolation). Let $\{(z_{1},f_{1}),\ldots,$ $(z_{\bar{d}},f_{\bar{d}})\}$ $\subseteq\mathbb{R}^{2}$ be a set of points whose values of $z_{d}$ are all distinct. Then, there exists a unique polynomial $f^{(\bar{d}-1)}$ of degree $\bar{d}-1$ that satisfies $f_{d}=f^{(\bar{d}-1)}(z_{d}),\forall d=1,\ldots,\bar{d}$. $\blacksquare$
Theorem 1 states that a minimum number of $d+1$ points equal to the degree of the polynomial are required to reconstruct the secret. This ensures information-theoretic security, meaning that even if an adversary obtains some shares, it is impossible to reconstruct the secret unless they have acquired the quorum number of shares.
Example 2 (Division of shares and secret reconstruction in SS). The procedures of Shamir’s SS [236], including the division of shares and the reconstruction of secrets, are given in Fig. 7.
Figure 7: An illustration of Shamir’s secret sharing, showing the division of shares and the secret reconstruction process.
Shamir’s SS can be executed in three steps: (1) Polynomial generation: A manager (secret holder) constructs a random polynomial $f(z)=s+c_{1}z+\cdots+c_{\bar{d}-1}z^{\bar{d}-1}$, where $s$ denotes the secret, the coefficients $c_{1},\ldots,c_{k-1}$ are randomly chosen from a uniform distribution in an integer field $\mathbb{E}\triangleq[0,e)$, where $e$ denotes a large prime number; (2) Division of shares: The manager computes the shares with a non-zero integer input and obtains the output, e.g., set $i=1,\ldots,n$ to retrieve Si $=(i,f(i)~{}(\bmod~{}e))$. Then, it distributes the share Si to the $i$th agent; (3) Secret reconstruction: Therefore, based on Theorem 1, at least $\bar{d}$ points are needed to reconstruct the polynomial and calculate the secret $s$. The efficient division of shares and reconstruction of secrets make SS highly suitable for privacy-preserving applications in multi-agent systems, including cryptographic key distribution, access control management in distributed systems [238], and the protection of sensitive data [239, 22].
SS has also shown potential in protecting privacy for controlling grid-edge energy resources for power system applications, including electricity theft detection [240] and vehicle-to-grid integration [241, 81]. Adopting SS, Nabil et al. [240] design a privacy-preserving detection scheme to identify electricity theft from malicious consumers. Only masked meter readings from consumers are collected and sent to the SO, ensuring privacy protection and preventing data leakage. In [81], an SS-based cooperative EV charging control protocol is developed to achieve overnight valley filling without compromising the privacy of EV owners’ charging profiles. The distributed protocol enjoys high computation efficiency and accuracy. In [241], a privacy-preserving communication protocol based on SS is proposed for vehicle-to-grid integration. The proposed protocol ensures that the existing battery charge level, the quantity of replenished energy, and the duration of EVs being plugged in remain undisclosed to aggregators.
In summary, as a threshold scheme based on polynomials and finite geometries, Shamir’s SS is well-suited for secure computation and key sharing in cryptographic applications among multiple stakeholders. Some potential future directions for SS-based methods: (1) Homomorphic secret sharing. Combining SS with homomorphic properties allows computations to be performed on the shared data without revealing the secret. It is worth investigating efficient homomorphic operations within the SS framework for privacy-preserving computations in scalable multi-agent systems. (2) Threshold cryptography in dynamic environments. Traditional SS requires a predefined number of participants. However, real-world power system applications often involve dynamic groups and changing environments. Research could focus on adapting SS to handle dynamic groups, where agents can join or leave without compromising the shared secret. (3) Power system applications. Future research could focus on how SS can be adapted in different power system communication topologies, such as federated, distributed, and decentralized mechanisms, to protect sensitive DER data.
@@ 4.2.2 Secret sharing-based methods
Table 4: Secret Sharing (SS)-based Scalable and Privacy-Preserving Methods.
Method Ref. Problem Structure Adversaries Key Features
@@
Table 4: (Continued)
Refers to active adversaries    Ref.:Reference    SS: Secret sharing
4.3 Other miscellaneous and emerging methods
4.3.1 State decomposition
4.3.2 Noise injection
4.3.3 Garbled circuit
Table 5: Miscellaneous and Emerging Scalable and Privacy-Preserving Methods.
Method Ref. Problem Structure Adversaries Key Features